help please!

If do=1 92=4 anty-an tant. Find the closed form of an.
Answer #1

Ques: > with autant tuz1 ... Cat?:? Given a = 1 and a₂=4 to anti- ant au- With q=1 and 9=4 We get, - ₂ = 9,+ 7 4=+1 6,19-1 an→ = Fla) - 9-9,2 = x( Flve) - )) + red flse) F(2)– 1 - 3 = x Flx) - x + xe? F(x) se? Else) x Flve) - F13) = -32-1 tz 7 (Flve)Let us denote by a= ust! b = 5-] 2 Then, ) F()= -(206+ (uta) (x-b) He shall decompose this fraction now So, het - (2x+1) = Aaeger, + (-) | F(x) - (ax+1) | (7) (-) t a 6 movi Sue We get LA ata | ㅋ 4 () C) () Similarly, 다 . > - - = 1. 5 1 2 o F(x) 네 +7 reso corte tra]* → FO) = bnt] Corro This, I we ano an= (-jht1 get anti tanto ntb We aga Since, a= √5 + 1 and, bo= 15 - 1 Ce